calculate the theroretical mean , standard dev, and variance simple size n=16 and process prob. = 0.10
Answer #1

Answer

Given :- sample size n = 16 and probability p = 0.10

Mean = n *p

= 16*0.10

= 1.6

And

Variance= n*p*q

where n = 100, p = 0.10 and q = 1-p = 1-0.10 = 0.90

this implies

variance = 16*0.10*0.90

= 1.44

and we know that standard deviation = sqrt{variance}

= sqrt{1.44}

= 1.2

So, mean = 1.6, variance = 1.44 and standard deviation = 1.2